OpenStudy (anonymous):

Find the radius of a sphere with a surface area of 1,236. (use 3.14 for pi) Answer 98.41 9.92 18.08 309

OpenStudy (anonymous):

surface area A=4pi r^2 =1236 r^2 =1236/(4 pi) r=sqrt(1236/(4 pi)) r=9.92 ans